OpenStudy (anonymous):

Solve x^2/3-3x^1/3-10=0

OpenStudy (anonymous):

\[x ^{2/3}-3x ^{1/3}-10=0\]

OpenStudy (anonymous):

\[put x ^{\frac{ 1 }{3 }}=y,x ^{\frac{ 2 }{3 }}=y ^{2}\]

OpenStudy (anonymous):

huh?

OpenStudy (shamil98):

let x^1/3 = y and x^2/3 = y^2 so y^2 - 3y - 10 = 0 (y-5)(y+2) = 0 sub back in (x^1/3 - 5)(x^1/3+2) = 0 make sense?

OpenStudy (anonymous):

either y-5=0,y=5, \[x ^{\frac{ 1 }{3 }}=5,cubing x=5^{3}=125\] similarly find for other.